OpenStudy (anonymous):

SOLVE ln(x-3)+ln(x+4)=1

OpenStudy (saifoo.khan):

\[\large \ln (x-3) + \ln (x+4) = \ln_e e\]

OpenStudy (anonymous):

uhm so what follows?

OpenStudy (saifoo.khan):

\[\ln(x-3)(x+4) = \ln_e e\] \[(x-3)(x+4) = e\]

OpenStudy (anonymous):

what else

OpenStudy (saifoo.khan):

Solve using a calc. Make a quadratic equation and so on.

OpenStudy (anonymous):

doesnt it hit on two places

OpenStudy (saifoo.khan):

Yes it will, one negative and one postitve.

OpenStudy (anonymous):

put the assignment tells me to round to the nearest 4th place if possible and it only has spot to answer which means my answer has to be one thing..

OpenStudy (anonymous):

so use the positive?

OpenStudy (saifoo.khan):

Yes!!

OpenStudy (saifoo.khan):

The value of 'x" can't be -ve.

OpenStudy (anonymous):

okay i see, can you help me through another one?

OpenStudy (anonymous):

x^210^x-2x10^x=3(10^x)